A goaltender has a save percentage of 0.920. This means that the probability of any single shot taken on the goaltender being a goal is 0.08. What would be the expected number of goals scored on this goaltender in a game where she faced 35 shots?
Answer #1

n = 35

p = 0.08

expexted number E(x) = n * p

E(x) = 35 * 0.08 = 2.8 goals

What would be the expected number of goals scored on this goaltender in a game where she faced 35 shots?

2.8 goals